# MUBen: Benchmarking the Uncertainty of Molecular Representation Models

### v3

@@ 5 Results and Analysis
Table 3 demonstrates a notable drop in prediction performance when backbone weights are fixed; and random splits outperform scaffold splits. This is consistent with intuition: if backbone models serve solely as feature extractors instead of a part of the trainable predictors, they are less expressive for downstream tasks. Additionally, in-distribution features tend to be more predictable. An interesting exception emerges in regression calibration error, where frozen backbones perform better and random splits score lower. Upon examining the predicted values, we note that predictions for random splits exhibit a sharper distribution, i.e., smaller $\hat{\sigma}$. This suggests that the models are more confident in regressing in-distribution data, aligning with our previous observation for Uni-Mol. Conversely, frozen backbones are less prone to overfitting due to their constrained expressiveness. This behavior underscores the original models’ capability to distinguish between in-distribution and OOD features and assign confidence scores with precision.
@@ 6 Conclusion, Limitation, and Future Works
We introduce MUBen, a benchmark that assesses an array of UQ methods across diverse categories, leveraging backbone models that incorporate different descriptors for multiple molecular property prediction tasks. Our results reveal that Deep Ensembles assures performance enhancement under all circumstances but with significant computation cost. Temperature scaling and MC Dropout are simple yet effective for classification, while BBP and SGLD are better suited for regression UQ. Among backbones, Uni-Mol, leveraging the expressiveness of 3D molecular conformations, is the most powerful but prone to overconfidence. Backbones with other descriptors are advantageous in different conditions and should be selected according to specific usage cases.
@@ Appendix A Dataset Details
MUBen primarily focuses on dataset split by molecule scaffolds, which separates training, validation and test features as much as possible. This creates out-of-distribution (OOD) test sets and signifies the model’s capability of processing patterns absent in the fine-tuning process. For all datasets, we use the conventional 8:1:1 training, validation, and test ratio. The raw datasets are available on the MoleculeNet website. [^1] https://moleculenet.org/ We take advantage of the pre-processed version shared by Zhou et al. (2023) and use the same splits as in their work. [^2] Repository at https://github.com/dptech-corp/Uni-Mol/tree/main/unimol and download at https://bioos-hermite-beijing.tos-cn-beijing.volces.com/unimol_data/finetune/molecular_property_prediction.tar.gz.
@@
To explore the impact of dataset splitting techniques, we conduct experiments on four classification datasets (BACE, BBBP, ClinTox, Tox21) and four regression datasets (ESOL, FreeSolv, Lipophilicity, QM7) using random splitting. Despite their limited sizes, these datasets effectively represent the overall performance of the models. Each dataset is divided into three separate 8:1:1 training-validation-test sets using random seeds of 00, $1$, and $2$. Each backbone-UQ combination is trained once per split. We then average the results over $3$ runs from the splits to determine the final metrics for each dataset. Deep Ensembles was the only exception, being trained $3$ times for each dataset split using different seeds, totaling $9$ training cycles per dataset. All other training settings were consistent with those in the scaffold-splitting experiments.
@@ Appendix B Backbone Models and Implementation Details
ChemBERTa is built with the HuggingFace Transformers library (Wolf et al., 2019), and its pre-trained parameters are shared through the Huggingface’s model hub. We retain the default model architecture and use the DeepChem/ChemBERTa-77M-MLM checkpoint for ChemBERTa’s weight initialization. [^3] Accessible at https://huggingface.co/DeepChem/ChemBERTa-77M-MLM. We employ the last-layer hidden state corresponding to the token [CLS] to represent the input SMILES sequence and attach the output heads specified by the tasks/UQ methods on top of it for property prediction. On all datasets, ChemBERTa is fine-tuned with a learning rate of $10^{-5}$, a batch size of $128$, and for $200$ epochs. A tolerance of $40$ epochs for early stopping is adopted.
@@
GROVER (Rong et al., 2020) is pre-trained on $11$M unlabeled molecules as 2D molecular graphs from ZINC15 (Sterling & Irwin, 2015) and ChEMBL (Gaulton et al., 2012). To enhance pre-training on molecular graphs, GROVER integrates the graph neural network (GNN)-based Message Passing Neural Networks (MPNN, Gilmer et al., 2017) into the Transformer encoder architectures. It also devises self-supervised objectives from different molecular structural granularities to better capture the rich structural and semantic information of molecules. Specifically, GROVER replaces the linear layer used for mapping the input query, key, and value vectors of each multi-head attention module into the head-dependent latent feature spaces in the Transformer encoder with a dynamic message passing network. This network aggregates the latent features from the $k$-hop neighboring nodes, where $k$ is a random integer drawn from a pre-defined Uniform of Gaussian. In other words, GROVER puts GNN layers before each head’s self-attention computation to better encode the graph structure of molecules. To facilitate learning, GROVER introduces training objectives including contextual property prediction, which predicts the statistical properties of masked subgraphs of arbitrary sizes, and graph-level motif prediction, which predicts the classes of the masked functional groups. GROVER uses a readout function to aggregate node features, which are then passed into linear layers for property prediction.
For implementation, we use the GROVER-base checkpoint for our model initialization. [^4] Repository available at https://github.com/tencent-ailab/grover, and model can be downloaded at https://ai.tencent.com/ailab/ml/ml-data/grover-models/pretrain/grover_base.tar.gz. We incorporate the “node-view” branch as discussed above and disregard the “edge-view” architecture detailed in the appendix of Rong et al. (2020), in accordance with the default settings in their GitHub repository. Under this configuration, the model generates $2$ sets of node embeddings (but no edge embeddings), one from the node hidden states and another from the edge hidden states (Rong et al., 2020). Each set of embeddings is passed into $2$ linear layers with GELU (Hendrycks & Gimpel, 2016) activation and a dropout ratio of $0.1$ post-readout layer, which simply averages the embeddings in the default configuration. Each set of embeddings corresponds to an individual output branch, predicting the properties independently. In line with the original implementation, we compute the loss for each branch individually during fine-tuning and apply a squared Euclidean distance regularization with a coefficient of $0.1$ between the two. During inference, we average the logits from the $2$ branches to generate the final output logits.
@@
Our codebase is developed atop the publicly accessible Uni-Mol repository, [^5] https://github.com/dptech-corp/Uni-Mol/tree/main. and their pre-trained checkpoint for molecular prediction serves as our model initialization. [^6] https://github.com/dptech-corp/Uni-Mol/releases/download/v0.1/mol_pre_no_h_220816.pt. During fine-tuning, Uni-Mol generates $10$ sets of 3D conformations for each molecule, supplemented with an additional 2D molecular graph. Thereafter Uni-Mol samples one from these $11$ molecular representations for each molecule at the beginning of every training epoch as the input feature. For inference, we average the logits from all $11$ representations to generate the final output. We utilize the conformations prepared by the Uni-Mol repository in our implementation.
@@
The Deep Neural Network (DNN) serves as a simple, randomly initialized baseline model designed to explore how heuristic descriptors like Morgan fingerprints (Morgan, 1965) or RDKit features perform for molecular property prediction. DNN enables us to compare the pre-trained models, which learn the molecular representation automatically through self-learning, with heuristic molecular features, which are constructed manually, and investigate whether or under what circumstances the heuristic features can achieve comparable results. For the descriptor, we adopt the approach of previous work (Wu et al., 2018; Yang et al., 2019; Rong et al., 2020) and extract $200$-dimensional molecule-level features using RDKit for each molecule, which are then used as DNN input. [^7] We use the “rdkit2dnormalized” descriptor in DescriptaStorus, available at https://github.com/bp-kelley/descriptastorus. The DNN consists of $8$ fully connected $128$-dimensional hidden layers with GELU activation and an intervening dropout ratio of $0.1$. We find no performance gain from deeper or wider DNNs and thus assume that our model is fully capable of harnessing the expressivity of RDKit features. The model is trained with a batch size of $256$ and a constant learning rate of $2\times 10^{-4}$ for $400$ epochs with an early stopping tolerance of $50$ epochs.
@@
To implement TorchMD-NET, we sourced the code and model checkpoint from (Zaidi et al., 2023). [^8] The repository can be accessed athttps://github.com/shehzaidi/pre-training-via-denoising. We made minor architectural adjustments, replacing their single-head output block with our adaptive multi-head output layers. Consequently, we omitted the denoising objective during the fine-tuning process due to compatibility concerns. Our fine-tuning regimen for TorchMD-NET entails a batch size of $128$ over $100$ epochs, adopting an early stopping mechanism with the patience of $40$ epochs. The learning rate peaks at $2\times 10^{-4}$, coupled with a linear scheduler with a $0.1$ warm-up ratio, and the model trains with a dropout ratio of $0.1$.
@@
We use the Pytorch Geometric (Fey & Lenssen, 2019) to realize GIN. Our implementation contains $5$ GIN layers with $128$ hidden units and $0.1$ dropout ratio. The model is trained with a batch size of $128$ for $200$ epochs with an early stopping tolerance of $50$ epochs, at a constant learning rate of $10^{-4}$.
@@ C.1 Method and Implementation Details
The focusing parameter $\gamma$ smoothly adjusts the rate at which easy examples are down-weighted. When $\gamma=0$, Focal Loss is equivalent to the cross-entropy loss. As $\gamma$ increases, the effect of the modulating factor increases likewise. In our implementation, we take advantage of the realization in the “torchvision” library and use their default hyper-parameters for all experiments. [^9] https://pytorch.org/vision/main/generated/torchvision.ops.sigmoid_focal_loss.html.
@@
$\bm{w}^{\rm MAP}=\operatorname*{arg\,max}_{\bm{w}}\log p(\bm{w}|\mathcal{D})=\operatorname*{arg\,max}_{\bm{w}}\left(\log p(\mathcal{D}|\bm{w})+\log p(\bm{w})\right),$ (2)
@@
However, the true posterior is generally intractable for neural networks and can only be approximated with variational inference $q(\bm{w}|\bm{\theta})$ (Kingma & Welling, 2014), where $\bm{\theta}$ are variational parameters. In our case, the variational parameters consist of the mean and standard deviation of multivariate Gaussian distribution $\bm{\theta}=\{\bm{\mu},\bm{\rho}\}$. The learning then involves finding the $\bm{\theta}$ that minimizes the Kullback-Leibler (KL) divergence between the true posterior and the variational distribution. The loss can be written as
$\mathcal{L}=\text{KL}\left[q(\bm{w}|\bm{\theta})||p(\bm{w})\right]-\mathbb{E}_{q(\bm{w}|\bm{\theta})}[\log p(\mathcal{D}|\bm{w})].$ (3)
@@
$\mathcal{L}_{i}\approx\log q(\bm{w}_{i}|\bm{\theta})-\log p(\bm{w}_{i})-\log p(\mathcal{D}|\bm{w}_{i}).$ (4)
@@ D.2 Uncertainty Quantification Metrics
Binary ECE puts the predicted probabilities $\{\hat{p}_{n}\}_{n=1}^{N}$ into $S$ (which is set to $S=15$ in our experiments) equal-width bins $B_{s}=\{n\in\{1,\dots,N\}\mid\hat{p}_{n}\in(\rho_{s},\rho_{s+1}]\}$ with $s\in\{1,\dots,S\}$. For binary classification, ECE is calculated as
$\begin{gathered}\text{ECE}=\sum_{s=1}^{S}\frac{|B_{s}|}{N}\left|\text{acc}(B_{s})-\text{conf}(B_{s})\right|;\\ \text{acc}(B_{s})=\frac{1}{|B_{s}|}\sum_{n\in B_{s}}y_{n};\quad\quad\text{conf}(B_{s})=\frac{1}{|B_{s}|}\sum_{n\in B_{s}}\hat{p}_{n}.\\ \end{gathered}$ (17)

### v4

@@ 5 Results and Analysis
Table 3 demonstrates a notable drop in prediction performance when backbone weights are fixed. Also, models evaluated on random splits outperform scaffold splits. This is consistent with intuition: if backbone models serve solely as feature extractors instead of a part of the trainable predictors, they are less expressive for downstream tasks. Additionally, in-distribution features tend to be more predictable. An interesting exception emerges in regression calibration error, where frozen backbones perform better and random splits score lower. Upon examining the predicted values, we note that predictions for random splits exhibit a sharper distribution, i.e., smaller $\hat{\sigma}$. This suggests that the models are more confident in regressing in-distribution data, aligning with our previous observation for Uni-Mol. Conversely, frozen backbones are less prone to overfitting due to their constrained expressiveness. This behavior underscores the original models’ capability to distinguish between in-distribution and OOD features and assign confidence scores with precision.
(a) RMSE $\downarrow$
(b) MAE $\downarrow$
(c) NLL $\downarrow$
(d) CE $\downarrow$
(e) RMSE $\downarrow$
(f) MAE $\downarrow$
(g) NLL $\downarrow$
(h) CE $\downarrow$
Figure 6: Performance of backbone and UQ methods according to the similarity between test data points and training scaffolds in the QM9 dataset. LABEL:sub@subfig:uqavg.rmseLABEL:sub@subfig:uqavg.ce illustrate the performance of backbone models averaged across all UQ methods, while LABEL:sub@subfig:backboneavg.rmseLABEL:sub@subfig:backboneavg.ce display the UQ performance with backbone averaged. The x-axis categorizes test data points into $5$ bins based on their average Tanimoto similarity to the training scaffolds, starting from the most similar (leftmost bin). The average Tanimoto similarities for bins 11115555 are $0.116$, $0.103$, $0.093$, $0.084$, and $0.066$, respectively. Lower values indicate better performance. The complete results are in Figure 10.
Impact of Training-Test Distribution Shift
Here we delve into the effects of data distribution on model performance within the QM9 dataset. We classify the test data into $5$ bins according to their average Tanimoto similarities to the training scaffolds, from the most to the least similar. Please refer to appendix A for more details on the dataset preparation. As depicted in Figure 6, there is a clear trend of worsening model performance, as evidenced by a nearly linear increase in RMSE and MAE, when the test data features gradually diverge from the training scaffolds. Conversely, the calibration error remains comparatively stable across these bins, suggesting that the estimated uncertainties are more robust to shifts in distribution, which reinforces our earlier findings regarding the behavior of backbone and UQ methods.
@@ 6 Conclusion, Limitation, and Future Works
We present MUBen, a benchmark designed to evaluate a variety of UQ methods across different categories, using backbone models that employ various descriptors for multiple molecular property prediction tasks. Our findings indicate that Deep Ensembles consistently enhance performance compared to the deterministic baseline, albeit at a substantial computational cost. For classification tasks, Temperature Scaling and MC Dropout are straightforward yet effective approaches. Conversely, for regression, BBP and SGLD appear more suitable in estimating uncertainty, although they may lead to a decrease in prediction accuracy, particularly with smaller backbone models. Among the different backbones, Uni-Mol stands out due to its effective utilization of 3D molecular conformations, which, while highly expressive, is also prone to overconfidence. Other backbone models, leveraging different descriptors, offer advantages under varying conditions and should be chosen based on the specific requirements of the use case.
@@ Appendix A Dataset Details
MUBen primarily utilizes a dataset split based on molecule scaffolds, effectively segregating training, validation, and test sets to maximize feature separation. This approach generates OOD test sets, crucial for assessing the model’s ability to handle patterns not encountered during the fine-tuning process. We adhere to the standard 8:1:1 ratio for training, validation, and test splits across all datasets. The raw datasets can be accessed on the MoleculeNet website. [^1] https://moleculenet.org/ Additionally, we employ the pre-processed versions provided by Zhou et al. (2023), using the identical dataset splits outlined in their study. [^2] https://github.com/dptech-corp/Uni-Mol/tree/main/unimol
@@
To investigate the impact of dataset splitting methods, we conducted experiments on $4$ classification datasets (BACE, BBBP, ClinTox, Tox21) and $4$ regression datasets (ESOL, FreeSolv, Lipophilicity, QM7) using random splitting. Despite their relatively small sizes, these datasets provide a reliable representation of model performance. Each dataset was divided into three separate training-validation-test sets with an 8:1:1 ratio, utilizing random seeds of 00, $1$, and $2$. We trained each backbone-UQ combination once per split and averaged the results across the three runs to establish the final metrics for each dataset. Deep Ensembles was the only exception; it underwent three training cycles per dataset split, using different seeds, resulting in a total of $9$ training cycles per dataset. All other training configurations remained consistent with those used in the scaffold-split experiments.
Binning Test Data by Similarity to Training Scaffolds
Our experimental analysis is conducted exclusively on the QM9 dataset due to its ample volume of test data, ensuring a statistically significant number of samples within each defined bin for reliable outcomes. Initially, the dataset is partitioned into training, validation, and test sets using scaffold splits. We then compute the average Tanimoto similarity between each test data point and the unique training scaffolds ($1096$ in total). Test data points are sorted and grouped into $5$ bins based on descending similarity scores, with the most similar points placed in the first bin. The average Tanimoto similarities for these $5$ bins are $0.116$, $0.103$, $0.093$, $0.084$, and $0.066$, respectively. During the experiments, each combination of backbone and UQ model is trained once on the training dataset and then tested separately across each of the bins. This process is repeated three times using random seeds 00, $1$, and $2$, with the results averaged across these runs. For the Deep Ensembles method, these three runs are integrated into a single ensemble, considered equivalent to a single run.
@@ Appendix B Backbone Models and Implementation Details
ChemBERTa is built with the HuggingFace Transformers library (Wolf et al., 2019), and its pre-trained parameters are shared through the Huggingface’s model hub. We retain the default model architecture and use the DeepChem/ChemBERTa-77M-MLM checkpoint for ChemBERTa’s weight initialization. [^3] https://huggingface.co/DeepChem/ChemBERTa-77M-MLM. We employ the last-layer hidden state corresponding to the token [CLS] to represent the input SMILES sequence and attach the output heads specified by the tasks/UQ methods on top of it for property prediction. On all datasets, ChemBERTa is fine-tuned with a learning rate of $10^{-5}$, a batch size of $128$, and for $200$ epochs. A tolerance of $40$ epochs for early stopping is adopted.
@@
GROVER (Rong et al., 2020) is pre-trained on $11$ million unlabeled molecules represented as 2D molecular graphs, sourced from ZINC15 (Sterling & Irwin, 2015) and ChEMBL (Gaulton et al., 2012). This model enhances pre-training by integrating Message Passing Neural Networks (MPNNs, Gilmer et al., 2017), a GNN-based method, into Transformer encoder architectures. GROVER employs self-supervised learning objectives that reflect different levels of molecular structural complexity to capture detailed structural and semantic information. Specifically, it replaces the linear layers that map the input query, key, and value vectors in each multi-head attention module of the Transformer encoder with a dynamic MPNN, which aggregates latent features from $k$-hop neighboring nodes, where $k$ is a random integer chosen from a predefined uniform or Gaussian distribution. In essence, GROVER integrates GNN layers before each head’s self-attention process to more effectively encode the molecular graph structure. To aid learning, GROVER introduces training objectives such as contextual property prediction, which estimates the statistical properties of masked subgraphs of varying sizes, and graph-level motif prediction, which identifies the classes of masked functional groups. A readout function is used to aggregate node features, which are subsequently processed by linear layers for property prediction.
For implementation, we use the GROVER-base checkpoint for our model initialization. [^4] https://github.com/tencent-ailab/grover We incorporate the “node-view” branch as discussed above and disregard the “edge-view” architecture detailed in the appendix of Rong et al. (2020), in accordance with the default settings in their GitHub repository. Under this configuration, the model generates $2$ sets of node embeddings (but no edge embeddings), one from the node hidden states and another from the edge hidden states (Rong et al., 2020). Each set of embeddings is passed into $2$ linear layers with GELU (Hendrycks & Gimpel, 2016) activation and a dropout ratio of $0.1$ post-readout layer, which simply averages the embeddings in the default configuration. Each set of embeddings corresponds to an individual output branch, predicting the properties independently. In line with the original implementation, we compute the loss for each branch individually during fine-tuning and apply a squared Euclidean distance regularization with a coefficient of $0.1$ between the two. During inference, we average the logits from the $2$ branches to generate the final output logits. Our implementation of GROVER does not incorporate external RDKit or Morgan fingerprints, diverging from the authors’ original implementation.
@@
Our codebase is developed atop the publicly accessible Uni-Mol repository,\@footnotemark and their pre-trained checkpoint for molecular prediction serves as our model initialization. During fine-tuning, Uni-Mol generates $10$ sets of 3D conformations for each molecule, supplemented with an additional 2D molecular graph. Thereafter Uni-Mol samples one from these $11$ molecular representations for each molecule at the beginning of every training epoch as the input feature. For inference, we average the logits from all $11$ representations to generate the final output. We utilize the conformations prepared by the Uni-Mol repository in our implementation.
@@
The Deep Neural Network (DNN) serves as a simple, randomly initialized baseline model designed to explore how heuristic descriptors like Morgan fingerprints (Morgan, 1965) or RDKit features perform for molecular property prediction. DNN enables us to compare the pre-trained models, which learn the molecular representation automatically through self-learning, with heuristic molecular features, which are constructed manually, and investigate whether or under what circumstances the heuristic features can achieve comparable results. For the descriptor, we adopt the approach of previous work (Wu et al., 2018; Yang et al., 2019; Rong et al., 2020) and extract $200$-dimensional molecule-level features using RDKit for each molecule, which are then used as DNN input. [^5] We use the rdkit2dnormalized descriptor in DescriptaStorus, available at https://github.com/bp-kelley/descriptastorus. The DNN consists of $8$ fully connected $128$-dimensional hidden layers with GELU activation and an intervening dropout ratio of $0.1$. We find no performance gain from deeper or wider DNNs and thus assume that our model is fully capable of harnessing the expressivity of RDKit features. The model is trained with a batch size of $256$ and a constant learning rate of $2\times 10^{-4}$ for $400$ epochs with an early stopping tolerance of $50$ epochs.
@@
To implement TorchMD-NET, we sourced the code and model checkpoint from (Zaidi et al., 2023). [^6] https://github.com/shehzaidi/pre-training-via-denoising We made minor architectural adjustments, replacing their single-head output block with our adaptive multi-head output layers. Consequently, we omitted the denoising objective during the fine-tuning process due to compatibility concerns. Our fine-tuning regimen for TorchMD-NET entails a batch size of $128$ over $100$ epochs, adopting an early stopping mechanism with the patience of $40$ epochs. The learning rate peaks at $2\times 10^{-4}$, coupled with a linear scheduler with a $0.1$ warm-up ratio, and the model trains with a dropout ratio of $0.1$.
Table 6: Impact of the number and dimensionality of hidden layers across various datasets. The selected hyper-parameters generally achieve reasonable performance, although they may not be optimal for all datasets.
Number of Layers (Dimension: 128) Layer Dimensionality (Number of Layers: 5)
Dataset: Metric 3 Layers 4 Layers 5 Layers 6 Layers 64 Dim 128 Dim 256 Dim 512 Dim
BBBP: ROC-AUC $\uparrow$ 0.5736 0.6233 0.6284 0.6223 0.5876 0.6284 0.6255 0.5977
Tox21: ROC-AUC $\uparrow$ 0.6644 0.6774 0.6832 0.6901 0.6584 0.6832 0.6924 0.6962
ESOL: RMSE $\downarrow$ 1.6741 1.664 1.6017 1.6383 1.9951 1.6017 1.6167 1.5868
FreeSolv: RMSE $\downarrow$ 2.3296 2.2348 2.1362 2.5788 4.0666 2.1362 2.8263 3.1005
@@
We use the Pytorch Geometric (Fey & Lenssen, 2019) to realize GIN. Our implementation contains $5$ GIN layers with $128$ hidden units and $0.1$ dropout ratio. A study of the hyperparameters of GIN is presented in Table 6. The model is trained with a batch size of $128$ for $200$ epochs with an early stopping tolerance of $50$ epochs, at a constant learning rate of $10^{-4}$.
@@ C.1 Method and Implementation Details
The focusing parameter $\gamma$ smoothly adjusts the rate at which easy examples are down-weighted. When $\gamma=0$, Focal Loss is equivalent to the cross-entropy loss. As $\gamma$ increases, the effect of the modulating factor increases likewise. In our implementation, we take advantage of the realization in the torchvision library and use their default hyperparameters for all experiments. [^7] https://pytorch.org/vision/main/generated/torchvision.ops.sigmoid_focal_loss.html.
@@
${\bm{W}}^{\rm MAP}=\operatorname*{arg\,max}_{{\bm{W}}}\log p({\bm{W}}|\mathbb{D})=\operatorname*{arg\,max}_{{\bm{W}}}\left(\log p(\mathbb{D}|{\bm{W}})+\log p({\bm{W}})\right),$ (2)
@@
However, the true posterior is generally intractable for neural networks and can only be approximated with variational inference $q({\bm{W}}|\bm{\theta})$ (Kingma & Welling, 2014), where $\bm{\theta}$ are variational parameters, which consist of the multivariate Gaussian mean and variance in our case. The learning then involves finding the $\bm{\theta}$ that minimizes the Kullback-Leibler (KL) divergence between the true posterior and the variational distribution. The loss can be written as
$\mathcal{L}=D_{\mathrm{KL}}\left(q({\bm{W}}|\bm{\theta})||p({\bm{W}})\right)-\mathbb{E}_{q({\bm{W}}|\bm{\theta})}[\log p(\mathbb{D}|{\bm{W}})].$ (3)
@@
$\mathcal{L}_{i}\approx\log q({\bm{W}}_{i}|\bm{\theta})-\log p({\bm{W}}_{i})-\log p(\mathbb{D}|{\bm{W}}_{i}).$ (4)
@@ D.2 Uncertainty Quantification Metrics
Binary ECE puts the predicted probabilities $\{\hat{p}_{n}\}_{n=1}^{N}$ into $S$ (which is set to $S=15$ in our experiments) equal-width bins $\mathbb{B}_{s}=\{n\in\{1,\dots,N\}\mid\hat{p}_{n}\in(\rho_{s},\rho_{s+1}]\}$ with $s\in\{1,\dots,S\}$. For binary classification, ECE is calculated as
$\begin{gathered}\text{ECE}=\sum_{s=1}^{S}\frac{|\mathbb{B}_{s}|}{N}\left|\text{acc}(\mathbb{B}_{s})-\text{conf}(\mathbb{B}_{s})\right|;\\ \text{acc}(\mathbb{B}_{s})=\frac{1}{|\mathbb{B}_{s}|}\sum_{n\in\mathbb{B}_{s}}y_{n};\quad\quad\text{conf}(\mathbb{B}_{s})=\frac{1}{|\mathbb{B}_{s}|}\sum_{n\in\mathbb{B}_{s}}\hat{p}_{n}.\\ \end{gathered}$ (17)